# Near instantaneous O(1) Analog Solver Circuit for Linear Symmetric Positive-Definite Systems

### v1

@@ III Preliminary Design: $n$-unknowns
In this design, the given symmetric system of equations $Ax=b$ is mapped to an equivalent system of resistors, where the unknown vector $x$ represents the voltages at the nodes. To simplify the system and avoid incorporation of active dynamic components, the right-hand-side vector $b$ is not directly assigned as the supplied current. Instead, a diagonal matrix $K_{s}$, with strictly nonnegative coefficients, is introduced. This matrix represents the conductance of the resistors that connect the supply nodes to the nodes corresponding to the unknowns. These resistors have conductance values denoted as $k_{si}$. The introduction of $K_{s}$ modifies the original system by subtracting this matrix from both sides of the equation, resulting in the following form:
@@ III-B Left-Hand-Side: $(A-K_{s})x$
For each resistive element in the matrix $(A-K_{s})$, the resistance values of the variable resistors $R_{pot1}$ and $R_{pot2}$ are set to be the absolute value of the inverse of the corresponding conductance calculated by Equations 5 and 6. Two additional control bits are associated with each resistive element:
@@ III-C System Testing
To evaluate the performance of the system, random symmetric matrices of varying sizes were generated using MATLAB R2022a [36] and then simulated in LTspice v17.1 [37]. The simulations were transient analysis simulations, where the supply voltage was modeled as a step function that transitions from zero to $x_{s}^{+}$ or $x_{s}^{-}$ at 1 $\mu s$. This setup was used to analyze the system settling time and the accuracy of the solution for the unknown vector $x$. The observed performance of the preliminary design across different subdomains of the linear system of equations $Ax=b$, is summarized as follows:
@@
    While these systems can be represented by the circuit, the system does not converge to the expected solution. In such cases, the voltage at the output node of at least one operational amplifier (opamp) in the negative resistance circuits reaches the amplifier maximum or minimum output voltage. This saturation indicates a positive feedback loop that results in circuit instability. This instability observation can be explained by using the analogy between resistive networks and spring networks as presented in [38]. In mechanical systems, a non-positive eigenvalue in the stiffness matrix (analogous to the conductance matrix) signifies instability, where small perturbations in displacement (analogous to voltage) result in unbounded motion.
@@
    Figure 9: Analysis of 400 random systems using the preliminary design. Plot (a) shows the maximum error between the converged solution and the theoretical values of each system. Plot (b) shows the maximum convergence time for each system
@@ IV Proposed Design: $2n$-unknowns
The primary motivation for this design is to reduce the number of components in the circuit, increase the range of problems that can be represented using strictly positive resistance, and reduce the settling time of the circuit. The design achieves this by transforming the original $n\times n$ system of equations into a new $2n\times 2n$ system that ensures that most resistive elements are positive, without introducing significant additional digital computations. The transformed system is expressed as:
$\begin{bmatrix}K_{A}&K_{B}\\ K_{B}&K_{A}\end{bmatrix}\begin{Bmatrix}x\\ -x\end{Bmatrix}=\begin{Bmatrix}b-K_{s}x\\ -b-K_{s}(-x)\end{Bmatrix},$ (14)
@@
where $D$ is a diagonal matrix with strictly non-negative elements. This transformation ensures that the off-diagonal terms of the sub-matrices $K_{A}$ and $K_{B}$ are negative, corresponding to positive resistive elements in the circuit. Importantly, no additional computations are required for these values, as they directly correspond to the off-diagonal terms of the original $A$ matrix, although assigned between different nodes. For positive resistive element between nodes $i$ and $j$ in the original system, the transformed system retains the same positive resistance between these nodes. Additionally, it introduces another positive resistive element with identical conductance between nodes $n+i$ and $n+j$. For negative resistive elements between nodes $i$ and $j$ in the original system, the transformed system replaces the negative resistance with two positive resistive elements of the same conductance. One resistor connects nodes $i$ and $n+j$, while the other connects nodes $j$ and $n+i$.
IV-A $D$-Matrix
@@
The choice of matrix $D$ plays a pivotal role in determining the domain of symmetric linear systems that will be positive definite in the transformed system and, therefore, solvable. In addition, it influences both the speed and the accuracy of the solution. A comparable, though less practical, transformation was proposed for support trees [34]. In that approach, $D$ is selected as the diagonal of the original matrix $A$, and $K_{s}$ is set to the zero matrix. However, such approach does not guarantee that a positive definite system will remain positive definite after transformation, thereby reducing the range of problems that can be addressed.
IV-A1 Stability Condition
@@
To ensure the stability of the system, it is essential to analyze the conditions under which the transformed system remains positive definite, as this is a key criterion for stability. Using the theorems for the determinants of block matrices [39], the eigenvalues $\lambda$ of the transformed system, for any choice of $D$, is derived. The eigenvalues are computed as follows:
$det\left(\begin{bmatrix}K_{A}-\lambda I_{n}&K_{B}\\ K_{B}&K_{A}-\lambda I_{n}\end{bmatrix}\right)=\\ det(K_{A}+K_{B}-\lambda I_{n})\ det(K_{A}-K_{B}-\lambda I_{n})$ (17)
@@
Thus, positive definiteness of the system is achieved if both matrices $(K_{A}-K_{B})$ and $(K_{A}+K_{B})$ are positive definite. These matrices can be expressed and simplified as follows:
$K_{A}-K_{B}$ $=A-K_{s},$ (18)
$K_{A}+K_{B}$ $=2D-|A|-K_{s}.$ (19)
@@
The condition expressed in Equation 18 requires that the original system $(A-K_{s})$ be positive definite. To preserve the positive definiteness of the transformed system, the matrix $D$ is chosen to ensure that $(K_{A}+K_{B})$ is diagonally dominant, making it positive definite. Since $(K_{A}+K_{B})$ can be decomposed as the sum of a diagonal matrix $D$ with strictly positive elements and a matrix $(-|A|-K_{s})$ with strictly non-positive elements, achieving diagonal dominance requires that the sum of each column of $(K_{A}+K_{B})$ be greater than zero. This requirement leads to the following condition for each element of the matrix $D$:
@@ IV-A2 Scaling
To optimize the design for the choice of $D$, its effect on both the accuracy and speed of the expected solution was studied. In this study, $D$ was selected as a scaled identity matrix, with the scaling determined as the maximum value of the sums of the columns of the absolute value of the conductance matrix $A$ as follows:
$D=\beta\max_{i}\left(\sum_{j=1}^{n}|A_{ji}|\right)I_{n},$ (21)
where $\beta$ is a scaling factor with $\beta\geq 0.5$ to satisfy the condition in Equation 20.
@@ IV-A3 Proposed $D$-Matrix
Figure 11: Proposed design in cross-point architecture
IV-A4 Cross-point Architecture
@@
One notable advantage of the proposed design is its compatibility with a cross-point architecture, as shown in Figure 11, the same architecture widely used for Matrix Vector Multiplication (MVM). In this setup, the rows and columns of the array represent the voltages of the unknown vector $x$, with each row directly connected to the corresponding column. In addition, two additional columns are introduced for the supply nodes $x_{s}^{+}$ and $x_{s}^{-}$, and an additional row is added for the ground node.
@@
The elements of the transformed matrix $K$ correspond to the conductance values between the various nodes in the unknown vector $x$. Theoretically, only half of the elements in the cross-point array are needed to represent the system due to symmetry. However, to maintain the matrix structure in the array for MVM operations, the off-diagonal terms of the submatrices $K_{A}$ and $K_{B}$ are divided by 2. These values are then assigned symmetrically to the elements $k_{ij}$ and $k_{ji}$, effectively splitting the original resistors into two parallel resistors.
@@
The diagonal elements of the transformed matrix $K$, whether assigned within the array or not, do not influence the design, as both ends of these resistors are connected to the same node, rendering them effectively nonexistent. Furthermore, the diagonals of the submatrix $K_{B}$ are deliberately set to zero in the array, as these elements can represent either positive or negative resistance. External element circuits are used to connect each node $i$ to node $n+i$, enabling flexible resistance representation while preserving the structure of the array for MVM operations.
@@
The elements of the supply matrix $K_{s}$ are assigned to the conductance values in the two additional columns, based on the sign of the corresponding element in the right-hand-side vector $b$. The conductance values are calculated using Equation 13. The additional row in the system includes the conductance values calculated as the sums of the columns of the transformed matrix $K$.
IV-B System Testing
To evaluate the performance of the proposed design, the same procedure as that used for the preliminary design was applied. Random symmetric matrices of varying sizes were generated using MATLAB and then simulated in LTspice v17.1. The simulations were transient analysis simulations, where the supply voltage was modeled as a step function that changes from zero to $x_{s}^{+}$ or $x_{s}^{-}$ at 1 $\mu s$. This setup was used to analyze the system’s settling time and the accuracy of the solution for the unknown vector $x$. By design, the domain of stable solutions for the proposed design corresponds to the domain of symmetric positive definite matrices.

### v2

@@ III Preliminary Design: $n$-unknowns
In this design, the given symmetric system of equations $Ax=b$ is mapped to an equivalent system of resistors, where the unknown vector $x$ represents the voltages at the nodes. To simplify the system and avoid the incorporation of active dynamic components, the right-hand-side vector $b$ is not directly assigned as the supplied current. Instead, a diagonal matrix $K_{s}$ with strictly nonnegative coefficients is introduced. This matrix represents the conductance of the resistors that connect the supply nodes to the nodes corresponding to the unknowns. These resistors have conductance values denoted as $k_{si}$. The introduction of $K_{s}$ modifies the original system by subtracting this matrix from both sides of the equation, resulting in the following form:
@@ III-B Left-Hand-Side: $(A-K_{s})x$
For each resistive element in the matrix $(A-K_{s})$, the resistance values of the variable resistors $R_{pot1}$ and $R_{pot2}$ are set to be the absolute value of the inverse of the corresponding conductance calculated by Equations (5) and (6). Two additional control bits are associated with each resistive element:
@@ III-C System Testing
To evaluate the performance of the system, random symmetric matrices of varying sizes were generated using MATLAB R2022a [44] and then simulated in LTspice v.17.1.8 [2]. The simulations were transient analysis simulations where the supply voltage was modeled as a step function that transitions from zero to $x_{s}^{+}$ or $x_{s}^{-}$ at 1 $\mu s$. This setup was used to analyze the system settling time and the accuracy of the solution for the unknown vector $x$. The observed performance of the preliminary design across different subdomains of the linear system of equations $Ax=b$, is summarized as follows:
@@
    While these systems can be represented by the circuit, the system does not converge to the expected solution. In such cases, the voltage at the output node of at least one operational amplifier (OA) in the negative resistance circuits reaches the maximum or minimum output voltage of the OA. This saturation indicates a positive feedback loop that results in circuit instability. This instability observation can be explained by using the analogy between resistive networks and spring networks as presented in [26]. In mechanical systems, a non-positive eigenvalue in the stiffness matrix (analogous to the conductance matrix) signifies instability, where small perturbations in displacement (analogous to voltage) result in unbounded response.
@@
    Figure 9: Analysis of 400 random dense systems (eigenvalues: [10 1000] $\mu S$) using the preliminary design. Plot (a) shows the maximum error between the converged solution and the theoretical values of each system. Plot (b) shows the maximum convergence time for each system.
@@ IV Proposed Design: $2n$-unknowns
The primary motivation for this design is to reduce the number of components in the circuit, increase the range of problems that can be represented using strictly positive resistance, and reduce the settling time of the circuit. The design achieves this by transforming the original $n\times n$ system of equations $Ax=b$ into a new $2n\times 2n$ system that ensures that most resistive elements are positive, without introducing significant additional digital computations. The transformed system $Ku=f$ is expressed as:
$\begin{bmatrix}K_{A}&K_{B}\\ K_{B}&K_{A}\end{bmatrix}\begin{Bmatrix}x\\ -x\end{Bmatrix}=\begin{Bmatrix}b-K_{s}x\\ -b-K_{s}(-x)\end{Bmatrix}$ (21)
@@
where $D$ is a diagonal matrix with strictly non-negative elements. This transformation ensures that the off-diagonal terms of the sub-matrices $K_{A}$ and $K_{B}$ are negative, corresponding to positive resistive elements in the circuit. Importantly, no additional computations are required for these values, as they directly correspond to the off-diagonal terms of the original $A$ matrix, although assigned between different nodes. For positive resistive element between nodes $i$ and $j$ in the original system, the transformed system retains the same positive resistance between these nodes. Additionally, it introduces another positive resistive element with identical conductance between nodes $n+i$ and $n+j$. For negative resistive elements between nodes $i$ and $j$ in the original system, the transformed system replaces the negative resistance with two positive resistive elements of the same conductance. One resistor connects nodes $i$ and $n+j$, while the other connects nodes $j$ and $n+i$. Thus, in the transformed $2n\times 2n$ system, the only off-diagonal terms that can possibly have a positive value (i.e. negative resistance), are the $n$ diagonal elements of $K_{B}$. This is a significant reduction from the possibly $n(n+1)/2$ negative resistances in the preliminary design shown in Figure 7.
IV-A Choice of $D$-Matrix
@@
The choice of matrix $D$ plays a pivotal role in determining the domain of symmetric linear systems that will be Positive Definite (PD) in the transformed system and, therefore, solvable. In addition, it influences both the speed and the accuracy of the solution. A comparable, though less practical, transformation was proposed for support trees [10]. In that approach, $D$ is selected as the diagonal of the original matrix $A$, and $K_{s}$ is set to the zero matrix. However, such approach does not guarantee that a positive definite system will remain positive definite after transformation, thereby reducing the range of problems that can be addressed.
IV-A1 Positive Definiteness Condition
@@
To ensure the stability of the system, it is essential to analyze the conditions under which the transformed system remains Positive Definite (PD), as this is a key criterion for stability. To find the eigenvalues of the transformed matrix $K$, we define an orthogonal matrix $Q$ that diagonalizes the block matrix $K$ as follows
$Q$ $=\frac{1}{\sqrt{2}}\begin{bmatrix}I&I\\ -I&I\end{bmatrix}$ (24)
$QKQ^{T}$ $=\begin{bmatrix}K_{A}+K_{B}&0\\ 0&K_{A}-K_{B}\end{bmatrix}$ (25)
The eigenvalues $\lambda$ of the transformed system, for any choice of $D$, are computed as follows:
$\det(K-\lambda I)=\det(K_{AB}-\lambda I_{n})\ \det(K_{A}-K_{B}-\lambda I_{n})$ (26)
@@
where $K_{AB}=K_{A}+K_{B}$. Thus, positive definiteness of the system is achieved if both matrices $(K_{A}-K_{B})$ and $K_{AB}$ are PD. These matrices can be expressed and simplified as follows:
$K_{A}-K_{B}$ $=A-K_{s}$ (27)
$K_{AB}=K_{A}+K_{B}$ $=2D-|A|-K_{s}$ (28)
@@
The condition expressed in Equation (27) requires that the original system $(A-K_{s})$ be PD. To preserve the positive definiteness of the transformed system, the matrix $D$ is chosen to ensure that $(K_{A}+K_{B})$ is diagonally dominant, making it PD. Since $(K_{A}+K_{B})$ can be decomposed as the sum of a diagonal matrix $D$ with strictly positive elements and a matrix $(-|A|-K_{s})$ with strictly non-positive elements, achieving diagonal dominance requires that the sum of each column of $(K_{A}+K_{B})$ be greater than zero. This requirement leads to the following condition for each element of the matrix $D$:
@@ IV-A2 Scaling
To optimize the design for the choice of $D$, the effect of the $D$ matrix on both the accuracy and speed of the expected solution was studied. In this study, $D$ was selected as a scaled identity matrix, with the scaling determined as the maximum value of the sums of the columns of the absolute value of the conductance matrix $A$ as follows:
$D=\beta\max_{i}\left(\sum_{j=1}^{n}|A_{ji}|\right)I_{n}$ (30)
where $\beta$ is a scaling factor with $\beta\geq 0.5$ to satisfy the condition in Equation (29).
@@ IV-A3 Proposed $D$-Matrix
Figure 11: Proposed design in cross-point architecture. Entries of the transformed $2n\times 2n$ matrix $K$ are directly assigned to the memristor array. The word lines (horizontal) and the bit lines (vertical) corresponding to the same entry of the unknown vector $x$ are connected. Two bit lines of memristors are added to represent the positive and negative entries of the right-hand side vector $b$. One word line is added to allow assigning conductance between each node and the ground. Peripheral circuitry representing the element circuit is also added between nodes $x_{i}$ and $x_{n+i}$ to account for any negative resistances corresponding to positive diagonals of $K_{B}$.
IV-B Cross-point Architecture
@@
A notable advantage of the proposed design is its compatibility with cross-point architecture, as shown in Figure 11, the same architecture that is widely used for Matrix Vector Multiplication (MVM). In this setup, the rows and columns of the array represent the voltages of the unknown vector $x$ (and x-x), with each row directly connected to the corresponding column. Two additional columns are introduced for the supply nodes $x_{s}^{+}$ and $x_{s}^{-}$, and an additional row is included for the ground node.
@@
Elements of the transformed matrix $K$ correspond to the conductance values between the various nodes in the unknown vector $x$. Theoretically, only half of the elements in the cross-point array are needed to represent the system due to symmetry. However, to maintain the matrix structure in the array for MVM operations, the off-diagonal terms of the submatrices $K_{A}$ and $K_{B}$ are divided by two. These values are then assigned symmetrically to the elements $k_{ij}$ and $k_{ji}$, effectively splitting the original resistors into two parallel resistors. In practical memristor-based implementations, double precision of the conductance values can be achieved by employing read–verify feedback programming across multiple arrays (typically no more than five) [38].
@@
The diagonal elements of the transformed matrix $K$, whether assigned within the array or not, do not influence the design, as both ends of these resistors are connected to the same node, rendering them effectively nonexistent. Furthermore, the diagonals of the submatrix $K_{B}$ are deliberately switched off in the array, as these elements can represent either positive or negative resistance. External element circuits are used to connect each node $i$ to node $n+i$, enabling flexible resistance representation while preserving the structure of the array for MVM operations.
@@
The elements of the supply matrix $K_{s}$ are assigned to the conductance values in the two additional columns, based on the sign of the corresponding element in the right-hand-side vector $b$. Positive elements $b^{+}$ are connected to $x_{s}^{+}$ whereas negative elements $b^{-}$ are connected to $x_{s}^{-}$. The conductance values are calculated using Equation (20). The additional row in the system includes the conductance values calculated as the sums of the columns of the transformed matrix $K$.
Once the circuit reaches steady state, the solution vector is obtained by reading the $n$ node voltages corresponding to the unknowns using ADCs connected to those nodes. For large arrays, the number of ADCs can be reduced by employing multiplexers along with the ADCs. Instead of reading all $n$ unknowns simultaneously, the solution voltages can be sampled serially in groups of $m$ nodes, where $m$ corresponds to the number of available ADCs.
To enable MVM operation on the proposed architecture, programmable switches can be introduced to reconfigure the peripheral circuitry between the element circuitry and the current sensing configuration required for MVM. Specifically, a set of switches selects between the peripheral element circuits and the operational amplifiers used for current sensing, while an additional set of switches on the word lines (WLs) temporarily disconnects the direct WL–bit line (BL) connections during MVM mode. The input vector for the MVM operation is applied by appropriately programming the $b^{+}$ and $b^{-}$ conductances using two-bit shifts, assuming symmetric supply voltages of $4\mathrm{V}$ and $4\mathrm{V}$.
IV-C Circuit Dynamics
A formal representation of the proposed design can be established by decomposing the transformed matrix $K$ into two matrices, $P$ and $N$, which correspond to the conductance matrices of the positive and negative resistance elements, respectively. The transformed system of equations $Ku=f$ can then be expressed as
(PN)u=f(P-N)u=f (33)
IV-C1 Governing Differential Equation
The proposed design guarantees that there is no more than one negative resistance circuit at each node $j$ in the system connected to node $n+j$. Let the diagonal matrix $R$ be the matrix of conductance values of the negative resistance circuits connected at each node. This matrix can be defined as the positive diagonals of the matrix $K_{B}$. The matrices $R$, $K_{N}$, $N$, and $P$ are computed as follows:
[…]
$K_{N}$ $=\begin{bmatrix}R&0\\ 0&R\end{bmatrix}$ (35)
$N$ $=\begin{bmatrix}R&-R\\ -R&R\end{bmatrix}$ (36)
$P$ $=K+N$ (37)
Applying Equation (17) for the transformed system and using the voltage relationship $u_{j}=-u_{n+j}$, the governing differential equation in vector form can be written as
$2\tau^{2}\frac{d^{2}u^{\prime}}{dt^{2}}+3\tau\frac{du^{\prime}}{dt}-2\tau\frac{du}{dt}=3u-u^{\prime}$ (38)
At any time $t$, the proposed design solves the resistive network that is composed from the positive conductance matrix $P$. The current supplied to the network is the summation of the current represented by the $f$-vector and the current supplied from the negative resistance circuit. The system in matrix form can be expressed as:
$Pu=f+K_{N}(u^{\prime}-u)$ (39)
Equation (39) can be rearranged as follows
$K_{N}u^{\prime}=Mu-f$ (40)
where $M=(P+K_{N})$. Taking the first and second derivatives of the equation results in the following:
$K_{N}\frac{du^{\prime}}{dt}$ $=M\frac{du}{dt}$ (41)
$K_{N}\frac{d^{2}u^{\prime}}{dt^{2}}$ $=M\frac{d^{2}u}{dt^{2}}$ (42)
Using the relationship $Nu=2K_{N}u$ that arises from the structure of the vector $u$, Equation (39) can be rewritten as:
$Ku-f=-K_{N}(3u-u^{\prime})$ (43)
Substituting Equations (41), (42), and (43) in Equation (38) gives:
$2\tau^{2}M\frac{d^{2}u}{dt^{2}}+3\tau M\frac{du}{dt}-2\tau K_{N}\frac{du}{dt}=-(Ku-f)$ (44)
To solve the second-order differential equation using state space, a new vector $w$ is defined as follows:
$w$ $=\tau\frac{du}{dt}$ (45)
$\frac{dw}{dt}$ $=\tau\frac{d^{2}u}{dt^{2}}$ (46)
Thus, the differential Equation (44) can be written as follows:
$2\tau\frac{dw}{dt}=M^{-1}(Ku-f)+(3I-2M^{-1}K_{N})w$ (47)
The matrix form of the complete differential equations:
[…]
where
$S=\begin{bmatrix}0&-2I\\ M^{-1}K&3I-2M^{-1}K_{N}\end{bmatrix}$ (49)
Using the finite difference, the governing equations become
$\begin{bmatrix}u(t+\Delta t)\\ w(t+\Delta t)\end{bmatrix}=\frac{\Delta t}{2\tau}\begin{bmatrix}0\\ M^{-1}f\end{bmatrix}+\left(I-\frac{\Delta t}{2\tau}S\right)\begin{bmatrix}u(t)\\ w(t)\end{bmatrix}$ (50)
$\begin{bmatrix}u(0)\\ w(0)\end{bmatrix}=\begin{bmatrix}M^{-1}f\\ 0\end{bmatrix}$ (51)
Figure 12: Differential Equation Validation. (a) Matrix AA of a random 16x16 linear system of equations that is composed of a mix of Diagonally Dominant ($D$) rows and non-DD rows. (b) Transient solution of the system using LTSpice simulation (color lines) and the finite difference algorithm of the derived differential equation (black dashed lines)
To validate the derived differential equation governing the circuit dynamics, 1000 random systems were generated and simulated in transient analysis using LTspice v.17.1.8. In these simulations, the OAs are modeled as single-pole models with a 10 MHz gain-bandwidth product. Across all tested systems, the transient response predicted by the differential equation matches within 0.2% the transient output voltages produced by the LTspice simulation. A representative example for a 16x16 system of equations containing a mixture of Diagonally Dominant columns (DD) and non-DD ones is shown in Figure 12.
IV-C2 Pole Analysis
To characterize the dynamic behavior of the circuit, the poles of the governing differential equation are analyzed. The dynamics of the full $2n$ system $K$ decouple into two independent $n$ subsystems, which arise from the diagonalization of the transformed matrix $K$ as shown in Equation (25). The first subsystem is associated with the matrix $K_{AB}$, whose eigenvectors lie in the vector space $\begin{bmatrix}I_{n}&I_{n}\end{bmatrix}^{T}$. As the unknown vector $u$ is orthogonal to this vector space, the convergence and stability of the solution process is unaffected by the dynamics of this subsystem. Nevertheless, it is essential that the $K_{AB}$ subsystem is not divergent, ensuring that perturbations do not grow and destabilize the circuit. The second subsystem corresponds to the system matrix $A$. This system governs the actual convergence behavior of the solver circuit.
The oscillation and stability conditions for the $K_{AB}$ system, given in the Appendix (Equations (84) and (87)), depend on the eigenvalues of the matrices $R$ and $K_{AB}^{-1}$. Since the matrix $R$ is a nonnegative diagonal matrix and the matrix $K_{AB}$ is diagonally dominant, both matrices have nonnegative eigenvalues. Consequently, the conditions in Equations (84) and (87) confirm that perturbations exciting the eigenvectors of the $K_{AB}$ subsystem neither oscillate nor grow, ensuring stability of these modes.
Similarly, the oscillation and stability conditions for the subsystem $A$, given in Equations (94) and (97), depend on the eigenvalues of the matrices $R$ and $A^{-1}$. If the matrix $A$ is SPD, these conditions indicate that the subsystem is overdamped and converges to the expected solution of the linear system of equations without oscillations. The convergence rate shown in Equation (101) depends on the largest eigenvalue of the matrix $A^{-1}R$. This eigenvalue can be bounded as follows:
$\lambda(A^{-1}R)\leq\frac{R_{max}}{\lambda_{min}(A)}$ (52)
where $R_{max}$ denotes the maximum value of the diagonal matrix $R$ and $\lambda_{min}(A)$ is the smallest eigenvalue of the matrix $A$. $R_{max}$ represents the maximum radius of the Gershgorin circles lying in the negative domain and may be considered as a measure of the maximum deviation from diagonal dominance.
IV-C3 Domain of Positive Resistors